OpenStudy (anonymous):

Radicle 125 over Radicle 4x^3

OpenStudy (anonymous):

\[\sqrt{125}/\sqrt{4x^3}\] \[\sqrt{125}\sqrt{4x^3}/4x^3\] \[\sqrt{500x^3}/4x^3\] \[10\sqrt{5x^3}/4x^3\]\[5\sqrt{5x^3}/2x^3\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I don't get the last two steps.